Express the following rational expressions in lowest terms:
1) x/x+xy ---------> My answer was xy
2) 3x^2+6x/3x^2+9x ---------> My answer was 2/3
3) x^2+2x+1/x+1 ---------> My answer was x+2x
4) x62+3x-18/2x-6
5) 2x62-4x-6/x-3
I tried the first 3 but I am not sure if I am right. The last two I don't understand. thank you for any help you can give me.

Answer by checkley75(3666)

1) X/X+XY=1+XY
2) (3X^2+6X)/(3X^2+9X)=3X(X+2)/3X(X+9) NOW CANCEL OUT THE 3X TERMS:
(X+2)/(X+9) ANSWER.
(3) X^2+2X+1)/(X+1)=(X+1)(X+1)/(X+1) CANCEL OUT THE (X+1) TERMS:
X+1 ANSWER.
(4) X^2+3X-18)/(2X-6)=(X-3)(X+6)/2(X-3) CANCEL OUT THE (X-3) TERMS:
(X+6)/2 ANSWER.
5) (2X^2-4X-6)/(X-3)=(2X+2)(X-3)/(X-3) CANCEL OUT THE (X-3) TERM:
2X+2 ANSWER.